OpenStudy (anonymous):

After a concert for 1,600 people, only 8 people said they would not go to see the band again. What percent of the people who went to the concert said they would not go to see the band again? 200 50% 8% 0.5% @Zeronknight @ZairaHernandez @ash_heather1 @Daniee_Bruhh @agent0smith @AYE_MY_PEEPS @agent0smith @aziaboo45

OpenStudy (anonymous):

8 is what percent of 1600? 8 is P% of 1600 Equation: Y = P% * X Solving the equation for P P% = Y/X P% = 8/1600 p = 0.005 Convert decimal to percent: P% = 0.005 * 100 = 0.5%